# Modeling wildfire dynamics through a physics-based approach incorporating fuel moisture and landscape heterogeneity

### v1

@@ 2.3 Modeling wind and topography
where we notice that $\beta<1$ in presence of fuel and $\beta=1$ when only air is present ($R_{f}=0$). Therefore, this coefficient can be regarded as a wind attenuation factor [27].
In [39], a similar expression for $\beta$ was derived without assuming thermal equilibrium between the gaseous and solid phases. This approach allowed for different values to be proposed inside and outside the flame. In the present work, we assume thermal equilibrium, i.e., $T_{f}=T_{a}=T$, and, as a result, $\beta$ does not depend on temperature. Furthermore, given the level of simplification in our model, we treat $\beta$ as a constant calibration parameter. However, variations arising from the effects of moisture, through $\bar{c}_{p,f}$, and the temporal evolution of $R_{f}$ might not be negligible. These aspects are further discussed in Sec. 4.4.
With the present formulation, a velocity field with non-zero divergence can appear. If solving the problem in conservative form, a correction is needed, compare Eq. (21). Alternatively, when solving the non-conservative version as we propose here no correction is required.
Approximating the effect of topography by means of an additional advective effect is a simple approach that allows to retain in the model’s solution the influence of topographic variations on wildfire propagation without introducing significant computational complexity [13, 50]. A different approach with a more meaningful physical basis would require modeling the effect of radiation globally, and not only locally through the Rosseland approximation [5]. Besides, we can find in the literature other approaches which provide a higher level of detail in the relation between $\mathbf{w}$ and the actual wind velocity 10 m above the ground [61].
For further investigations, we note that certain topographic influences are similar to an artificial wind velocity under this formulation. For example, a constant topographic slope transfers with the parameter $\gamma$ into a constant artificial wind velocity. If the fire is moving upslope, the artificial wind velocity is positive in this direction. If the fire moves downslope, the artificial wind velocity is negative. Experiments indicate that the linear assumption is acceptable for topographic slopes up to $30^{\circ}$ [56]. Any further investigations that are carried out for a wind velocity can be transferred easily to a topography with a constant slope.
@@ 2.4 Modeling local radiation
In the model herein presented, we only consider the local effect of radiation assuming that the medium is optically thick [4, 27, 65]. The diffusion coefficient in Eq. (23) includes some linear diffusion by conduction and a nonlinear term due to radiation that results from the Rosseland approximation. There is previous work where only the linear diffusion by conduction was considered [28, 42, 64]. For calibrated parameters, the numerical simulations show a small influence of the nonlinear diffusion by radiation, see [13, 65].
In this work, we investigate the influence of the radiation regarding two aspects: the overall influence of radiation and the dependency of the radiation term. The radiation depends on some fixed physical parameters, on the temperature $T$ and on the optical path length $\delta$. The optical path length is a physical value that is difficult to estimate. The challenges are caused by the unknown structure of the vegetation and by the reduction of the spatial dimensions from three to two. Therefore, we will keep $\delta$ constant in the model following previous literature [4, 27].
The key aspect when considering local radiation resides in the strong nonlinearity of the diffusion term. The dependency with the cubed temperature makes the diffusion mechanism very sensitive to this problem variable. In the parametric analyses in the results section, we will show that this dependency is required to reproduce physically consistent results.
@@ 2.5 Modeling combustion
that triggers the combustion when the temperature is higher than the pyrolysis temperature $T_{pc}$. In what follows, we assume that $T_{ac}=T_{pc}$.
Originally, molecular chemical processes motivated the Arrhenius law. In the wildfire model, the variables are acting on a much larger length scale, far away from the molecular scale. This motivates the approximation of the nonlinear Arrhenius law by a constant function as [5]
@@
The switch $s(T)$ in Eq. (36) for activating the combustion process above the ignition temperature is included for both approaches. Previous studies, see [34], show that the process of switching on is more important than the switching off. One can argue that the combustion processes continue even for temperature values below the ignition temperature once the fire was burning in this space. On the other hand, the switching off after the burning process affects only the cooling process and the remaining biomass fraction. It does not affect the maximal fire temperature nor the rate of spread of the fire. Our focus is more on the rate of spread. Therefore, we follow the approach in [42] and model the switch for the ignition and the cooling process.
We compare the two approaches for the combustion function $\Psi$ in Sec. 4.1 by means of numerical simulations.
@@ 2.6 Modeling fuel moisture
The bulk specific heat $\bar{c}_{p0}$ in Eq. (8) is a function of the specific heat of the fuel $\bar{c}_{p,f}$. In general, we can write the specific heat as a function of the temperature. However, it can also be interesting to consider its dependency on the biomass fraction in order to model the effect of fuel moisture evaporation and degradation of the biomass, i.e. $\bar{c}_{p,f}=\bar{c}_{p,f}(T,Y)$.
In this section, we will use the apparent calorific capacity method [55] to model the effect of fuel moisture. This method models the effect of thermal phase changes by considering an apparent (or effective) heat capacity, usually defined as a piecewise constant function. Thus, we will define $\bar{c}_{p,f}$ as a piecewise constant function featuring an apparent specific heat to account for the effect of fuel moisture.
We introduce four different approaches for the definition of $\bar{c}_{p,f}$. The first approach is referred to as NM (neglects moisture) and considers the specific heat of dry wood. The second approach is called the simple two-stage moisture model (S2M) and features a piecewise definition of the apparent specific heat with two stages in the heating process. The third approach is the three-stage moisture model (S3M) and considers three stages, being able to represent evaporation at a quasi-constant temperature. The fourth approach is called the complete two-stage moisture model (C2M), which is an extension of the S2M approach to distinguish between green and dead wood as the fuel of the fire.
@@ 2.6.1 Neglecting moisture (NM)
If we assume zero fuel moisture, we simply consider a constant specific heat for the fuel, i.e. the specific heat of the dry fuel,
@@ 2.6.2 Simple two-stage moisture model (S2M)
As a first approach to model fuel moisture, we define the moisture content $M$ as the ratio between the mass of water, $m_{w}$, and the mass of dry wood, $m_{0}$, as
@@
For consistency with Eq. (27) and to agree with literature standards [3, 43], we assume a dry basis. The increment of specific enthalpy when heating from the ambient temperature, $T_{\infty}$, to the evaporation temperature of water, $T_{w}$, before evaporation (sensible heating) is defined as
@@
where $\rho_{f_{0}}$ is the density of the dry fuel, $c_{p,f_{0}}$ is the specific heat of the dry fuel and $c_{w}$ is the specific heat of water. For the sake of simplicity, we assume that $\rho_{f_{0}}\approx\bar{\rho}_{f}$, which was defined in Sec. 2.1 as an input parameter. Accordingly, the increment of specific enthalpy during evaporation (latent heating) is defined as
@@ 2.6.3 Simple three-stage moisture model (S3M)
We now define an approximate enthalpy change for the sensible heating process
@@ 2.6.4 Complete two-stage moisture model (C2M)
The amount of water in the wood affects the process of heating the fuel from ambient temperature ($T_{\infty}$) to the pyrolysis temperature ($T_{pc}$). In the S2M model, we consider the scenario of fuel accumulating water homogeneously. Although this simplification can be useful in certain situations, in a more realistic approach the moisture content depends on the type of fuel which, for simplicity, will be referred here as wood. In this regard, we distinguish between two types of wood: dead and live (or green) wood. The second type refers to the trees or bushes which are still alive.
@@
Another important concept for dead wood is the equilibrium moisture content (EMC), which is defined as the moisture content at which the wood is neither gaining nor losing water. If the environmental conditions change, wood does not reach immediately the EMC. In fact, the rate at which dead wood exchanges moisture with the atmosphere to reach EMC depends crucially on the size of the sample. Dead fuel is separated into fine fuels –leaves, bark and twigs with a diameter lower than 6 mm– and coarse fuels—larger twigs and logs [31]. The US National Fire Danger Rating System [12] classifies fuel sizes according to the time it takes to reach equilibrium with the atmosphere. Fine fuels are within the 1-hour fuel class, whereas coarse fuels are within the 10-, 100- and 1000-hour classes. However, in this study we will assume no variation in time of the relative humidity and thus consider that the EMC is equal to $M_{d}$; for a more complete approach representing time variations, see for example [38]. In Tab. 1 we show values EMC, hereafter $M_{d}$, at $T=300K$.
@@
Eq. (56) is represented in Fig. 2, where it is shown the dependence of the moisture content $M_{d}$ on the temperature and on the relative humidity.
As in Eq. (43), we find the enthalpy balance
@@
where $c_{p}$ is the specific heat of the mixture of moist fuels. Then, we compute the effective specific heat adapted to the initial conditions of environment and wood,
$c_{p,\mathrm{eff}}=\frac{\Delta h_{p}}{T_{pc}-T_{\infty}},$ (59)
and take the approach
@@
which is depicted in Fig. 1(a) as a straight blue dashed line.
The heat capacity of wood depends on the temperature and moisture content of the wood, but it is practically independent of density or species. In order to compute $c_{p}$ in Eq. (58), referred to the mixture of wood from the value of the moisture content $M_{g}$ and $M_{d}$, we use the expressions
@@
where $A_{i}$ is a correction factor required below fiber saturation which accounts for the water accumulated at the cell walls of the wood where, as said above, it forms a bound state. An empirical formula for the value of $A_{i}$ depending on the temperature and moisture content is [43]
@@
to be activated below fiber saturation ($M_{i}<0.3$) where the correction factor $A_{i}$ needs to be taken into account. Above fiber saturation ($M_{i}\geq 0.3$) we can simply use a mixture rule to estimate the specific heat. For green wood we usually have $M_{g}>0.3$ [43], therefore the switch is zero, $s^{\prime}(M_{g})=0$.
@@
Figure 2: Moisture content $M_{d}$ depending on the temperature and the relative humidity, following Eq. (56).
@@ 2.7 Model parameters
This section summarizes the parameters required in the model. Tab. 2 includes environmental (ambient forcing) and fuel-dependent parameters. Although these parameters may exhibit spatial heterogeneity, the table presents their default values. Tab. 3 lists additional model parameters which, despite having a physical basis, are treated as calibration parameters and should not vary with spatial position. Among these are the optical path length, wind and slope correction factors, and the ambient cooling coefficient. Note that the latter does not correspond to a physically realistic value, as it also accounts for heat dissipation by radiation in the vertical direction, which is not explicitly represented in the model. Finally, Tab. 4 contains other parameters and physical constants, which are fixed.
@@
Parameter Symbol Default value Units
Ground elevation $Z$ m
Wind velocity $\mathbf{w}$ m·s-1
Surface coverage $SC$ 1.0 -
Initial fuel volume ratio $R_{f,0}$ 0.01 -
Fuel density $\bar{\rho}_{f}$ 400 kg·m-3
Initial fuel load $W_{f,0}$ 4 kg·m-2
Specific heat of dry fuel $c_{p,f0}$ 1.0 kJ·kg-1·K-1
Fuel pyrolysis temperature $T_{pc}$ 600 K
Fuel combustion heat $\mathcal{H}$ 4000 kJ·kg-1
Reaction rate $A_{c}$ 0.01 s-1
Pre-exponential factor $A_{a}$ 0.0173 s-1
Fuel moisture content§ $M$ 0.1 -
Live (green) fuel moisture $M_{g}$ 0.1 -
Live to dead fuel ratio $r$ 1.0 -
Relative humidity $\phi$ 0.3 -
@@
Parameter Symbol Default value Units
Thermal conductivity $k_{c}$ 0.0001 kW·m-1· K-1
Optical path length $\delta$ 1.0 m
Ambient cooling coefficient $\hat{\alpha}$ 0.05 kW·m-2·K-1
Wind correction coefficient $\beta$ 0.02 -
Slope correction coefficient $\gamma$ 0.04 -
@@ 3.1 Finite Volume spatial discretization
${\mathcal{L}(\mathbf{{U}}_{ij})=-\frac{(\mathbf{A}\Delta_{x}\mathbf{{U}})_{i,j}}{\Delta x}-\frac{(\mathbf{B}\Delta_{y}\mathbf{{U}})_{i,j}}{\Delta y}+\mathbf{C}_{ij}\left(\frac{\mathbf{{F}}_{i+1/2,j}-\mathbf{{F}}_{i-1/2,j}}{\Delta x}+\frac{\mathbf{{G}}_{i,j+1/2}-\mathbf{{G}}_{i,j-1/2}}{\Delta y}+\bar{\textbf{S}}_{ij}\right)\,},$ (68)
@@
The non-conservative products are only non-zero for the first equation and are computed using a high-order (7-th order) WENO reconstruction [53, 42]. The diffusive fluxes are only non-zero for the first equation, that is
@@
where $\widetilde{k}=0.5(k_{i+1}+k_{i})$, so it is the mean of the nonlinear diffusion in Eq. (23).

### v2

@@ 2.3 Modeling wind and topography
with $\beta<1$ in presence of fuel and $\beta=1$ when only air is present ($R_{f}=0$). Therefore, this coefficient can be regarded as a wind attenuation factor [29].
In [42], a similar expression for $\beta$ was derived without assuming thermal equilibrium between the gaseous and solid phases. This approach allowed for different values to be proposed inside and outside the flame. In the present work, thermal equilibrium, i.e., $T_{f}=T_{a}=T$, is assumed and, as a result, $\beta$ does not depend on temperature. Furthermore, given the level of simplification in the model, $\beta$ is treated as a constant calibration parameter. However, variations arising from the effects of moisture, through $\bar{c}_{p,f}$, and the temporal evolution of $R_{f}$ might not be negligible. These aspects are further discussed in Sec. 4.4.
With the present formulation, a velocity field with non-zero divergence can appear. If solving the problem in conservative form, a correction is needed [40]. Alternatively, when solving the non-conservative version as proposed no correction is required here.
Approximating the effect of topography by means of an additional advective effect is a simple approach that allows to retain in the model’s solution the influence of topographic variations on wildfire propagation without introducing significant computational complexity [13, 53]. A different approach with a more meaningful physical basis would require modeling the effect of radiation globally, and not only locally through the Rosseland approximation [5]. Besides, there are other approaches in the literature which provide a higher level of detail in the relation between $\mathbf{w}$ and the actual wind velocity 10 m above the ground [64].
For further investigations, note that certain topographic influences are similar to an artificial wind velocity in this formulation. For example, a constant topographic slope transfers with the parameter $\gamma$ into a constant artificial wind velocity. If the fire is moving upslope, the artificial wind velocity is positive in this direction. If the fire moves downslope, the artificial wind velocity is negative. Experiments indicate that the linear assumption is acceptable for topographic slopes up to $30^{\circ}$ [59]. Any further investigations that are carried out for a wind velocity can be transferred easily to a topography with a constant slope.
@@ 2.4 Modeling local radiation
In the model herein presented, only the local effect of radiation is considered, assuming that the medium is optically thick [4, 29, 68]. The diffusion coefficient in Eq. (28) includes some linear diffusion by conduction and a non-linear term due to radiation that results from the Rosseland approximation. There is previous work where only the linear diffusion by conduction was considered [30, 45, 67]. For calibrated parameters, the numerical simulations show a small influence of the non-linear diffusion by radiation, see [13, 68].
In this work, the influence of the radiation is investigated regarding two aspects: the overall influence of radiation and the dependency of the radiation term. The radiation depends on some fixed physical parameters, on the temperature $T$ and on the optical path length $\delta$. The optical path length is a physical value that is difficult to estimate. The challenges are caused by the unknown structure of the vegetation and by the reduction of the spatial dimensions from three to two. Therefore, $\delta$ is kept constant in the model following previous literature [4, 29].
The key aspect when considering local radiation resides in the strong non-linearity of the diffusion term. The dependency with the cubed temperature makes the diffusion mechanism very sensitive to this problem variable. The parametric analyses in the results section will show that this dependency is required to reproduce physically consistent results.
@@ 2.5 Modeling combustion
that triggers the combustion when the temperature is higher than the pyrolysis temperature $T_{pc}$. From this point onwards, it will be assumed that $T_{ac}=T_{pc}$. Assume also that pyrolysis and combustion occur sequentially, with gas oxidation beginning immediately after fuel devolatilization starts, thereby generating net exothermic heat. This allows $T_{pc}$ to be considered as the threshold for combustion.
Originally, molecular chemical processes motivated the Arrhenius law. In the wildfire model, the variables are acting on a much larger length scale, far away from the molecular scale. This motivates the approximation of the non-linear Arrhenius law by a constant function as [5]
@@
The switch $s(T)$ in Eq. (41) for activating the combustion process above the ignition temperature is included for both approaches. Previous studies, see [37], show that the process of switching on is more important than the switching off. One can argue that the combustion processes continue even for temperature values below the ignition temperature once the fire was burning in this space. On the other hand, the switching off after the burning process affects only the cooling process and the remaining biomass fraction. It does not affect the maximal fire temperature nor the rate of spread of the fire. This paper focuses more on the rate of spread. Therefore, following the approach in [45], the switch is modeled for the ignition and the cooling process.
Sec. 4.1 compares the two approaches for the combustion function $\Psi$ by means of numerical simulations.
@@ 2.6 Modeling fuel moisture
The bulk specific heat $\bar{c}_{p0}$ in Eq. (8) depends on the specific heat of the fuel $\bar{c}_{p,f}$. In general, the specific heat of the fuel is a function of the temperature. However, it can also be interesting to consider its dependency on the biomass fraction in order to model the effect of fuel moisture evaporation and degradation of the biomass, i.e. $\bar{c}_{p,f}=\bar{c}_{p,f}(T,Y)$.
In this section, the apparent calorific capacity method [58] is used to model the effect of fuel moisture. This method models the effect of thermal phase changes by considering an apparent (or effective) heat capacity, usually defined as a piecewise constant function. Thus, the specific heat $\bar{c}_{p,f}$ will be defined as a piecewise constant function featuring an apparent specific heat to account for the effect of fuel moisture.
Four different approaches for the definition of $\bar{c}_{p,f}$ are introduced. The first approach is referred to as NM (neglects moisture) and considers the specific heat of dry wood. The second approach is called the simple two-stage moisture model (S2M) and features a piecewise definition of the apparent specific heat with two stages in the heating process. The third approach is the three-stage moisture model (S3M) and considers three stages, being able to represent evaporation at a quasi-constant temperature. The fourth approach is called the complete two-stage moisture model (C2M), which is an extension of the S2M approach to distinguish between green and dead wood as the fuel of the fire.
@@ 2.6.1 Neglecting moisture (NM)
Assuming zero fuel moisture, the specific heat for the fuel can be considered constant, i.e. identical to the specific heat of the dry fuel,
@@ 2.6.2 Simple two-stage moisture model (S2M)
As a first approach to model fuel moisture, the moisture content $M$ is defined as the ratio between the mass of water, $m_{w}$, and the mass of dry wood, $m_{0}$, as
@@
For consistency with Eq. (32) and to agree with literature standards [3, 46], a dry basis is assumed. The increment of specific enthalpy when heating from the ambient temperature, $T_{\infty}$, to the evaporation temperature of water, $T_{w}$, before evaporation (sensible heating) is defined as
@@
where $\rho_{f_{0}}$ is the density of the dry fuel, $c_{p,f_{0}}$ is the specific heat of the dry fuel and $c_{w}$ is the specific heat of water. For simplicity, assume $\rho_{f_{0}}\approx\bar{\rho}_{f}$, which was defined in Sec. 2.1 as an input parameter. Accordingly, the increment of specific enthalpy during evaporation (latent heating) is defined as
@@ 2.6.3 Simple three-stage moisture model (S3M)
An approximate enthalpy change for the sensible heating process is then defined by
@@ 2.6.4 Complete two-stage moisture model (C2M)
The amount of water in the wood affects the process of heating the fuel from ambient temperature ($T_{\infty}$) to the pyrolysis temperature ($T_{pc}$). In the S2M model, the fuel accumulates water homogeneously. Although this simplification can be useful in certain situations, in a more realistic approach the moisture content depends on the type of fuel which, for simplicity, will be referred here as wood. In this regard, two types of wood are distinguished: dead and live (or green) wood. The second type refers to the trees or bushes which are still alive.
@@
Another important concept for dead wood is the equilibrium moisture content (EMC), which is defined as the moisture content at which the wood is neither gaining nor losing water. If the environmental conditions change, wood does not reach immediately the EMC. In fact, the rate at which dead wood exchanges moisture with the atmosphere to reach EMC depends crucially on the size of the sample. Dead fuel is separated into fine fuels –leaves, bark and twigs with a diameter lower than 6 mm– and coarse fuels—larger twigs and logs [33]. The US National Fire Danger Rating System [12] classifies fuel sizes according to the time it takes to reach equilibrium with the atmosphere. Fine fuels are within the 1-hour fuel class, whereas coarse fuels are within the 10-, 100- and 1000-hour classes. However, in this study no variation in time of the relative humidity is assumed and thus $M_{d}$ can be approximated by the EMC; for a more complete approach representing time variations, see for example [41]. Tab. 1 gives values of EMC, hereafter $M_{d}$, at $T=300K$.
@@
As in Eq. (48), the enthalpy balance is
@@
where $c_{p}$ is the specific heat of the mixture of moist fuels. Then, the effective specific heat adapted to the initial conditions of environment and wood is
$c_{p,\mathrm{eff}}=\frac{\Delta h}{T_{pc}-T_{\infty}},$ (64)
and the approach
@@
is displayed as a straight blue dashed line in Fig. 1(a).
The heat capacity of wood depends on the temperature and moisture content of the wood, but it is practically independent of density or species. In order to compute $c_{p}$ in Eq. (63), referred to the mixture of wood from the value of the moisture content $M_{g}$ and $M_{d}$, the expressions
@@
are used, where $A_{i}$ is a correction factor required below fiber saturation which accounts for the water accumulated at the cell walls of the wood where, as said above, it forms a bound state. An empirical formula for the value of $A_{i}$ depending on the temperature and moisture content is [46]
@@
to be activated below fiber saturation ($M_{i}<0.3$) where the correction factor $A_{i}$ needs to be taken into account. Above fiber saturation ($M_{i}\geq 0.3$) a mixture rule is used to estimate the specific heat. For green wood, it is known that is $M_{g}>0.3$ [46], therefore the switch is zero, $s^{\prime}(M_{g})=0$.
@@ 2.7 Model parameters
This section summarizes the parameters required in the model. Tab. 2 includes environmental (ambient forcing) and fuel-dependent parameters. Although these parameters may exhibit spatial heterogeneity, the table presents their default values. Tab. 3 lists additional model parameters which, despite having a physical basis, are treated as calibration parameters and should not vary with spatial position. Among these are the optical path length, wind and slope correction factors, and the ambient cooling coefficient. Note that the latter does not correspond to a physically realistic value, as it also accounts for the energy loss owing to the mass escaping through the top of the domain, as well as for heat dissipation by radiation in the vertical direction, which is not explicitly represented in the model. Finally, Tab. 4 contains other parameters and physical constants, which are fixed.
The default parameter values in the tables are based on previous studies and preliminary calibration. Rounded values are used, consistent with the order of magnitude found in the reference sources provided in the tables.
@@
Parameter Symbol Default value Units Ref.
Ground elevation $Z$ m -
Wind velocity $\mathbf{w}$ m·s-1 -
Surface coverage $SC$ 1.0 - -
Initial fuel volume ratio $R_{f,0}$ 0.01 - [55, 32]
Fuel density $\bar{\rho}_{f}$ 400 kg·m-3 [35]
Initial fuel load $W_{f,0}$ 4 kg·m-2 [32]
Specific heat of dry fuel $c_{p,f0}$ 1.0 kJ·kg-1·K-1 [35]
Fuel pyrolysis temperature $T_{pc}$ 600 K [29]
Fuel combustion heat $\mathcal{H}$ 4000 kJ·kg-1 -
Reaction rate $A_{c}$ 0.01 s-1 -
Pre-exponential factor $A_{a}$ 0.0173 s-1 -
Fuel moisture content§ $M$ 0.1 - [46]
Live (green) fuel moisture $M_{g}$ 0.9 - [46]
Live to dead fuel ratio $r$ 1.0 - -
Relative humidity $\phi$ 0.3 - -
@@
Parameter Symbol Default value Units Ref.
Thermal conductivity $k_{c}$ 0.0001 kW·m-1· K-1 [29]
Optical path length $\delta$ 1.0 m [29, 32]
Ambient cooling coefficient $\hat{\alpha}$ 0.05 kW·m-2·K-1 [32]
Wind correction coefficient $\beta$ 0.02 - -
Slope correction coefficient $\gamma$ 0.04 - -
@@ 3.1 Finite Volume spatial discretization
$\mathcal{L}(\mathbf{{U}}_{ij})=-\frac{(\mathbf{A}\Delta_{x}\mathbf{{U}})_{i,j}}{\Delta x}-\frac{(\mathbf{B}\Delta_{y}\mathbf{{U}})_{i,j}}{\Delta y}+\mathbf{C}_{ij}\left(\frac{\mathbf{{F}}_{i+1/2,j}-\mathbf{{F}}_{i-1/2,j}}{\Delta x}+\frac{\mathbf{{G}}_{i,j+1/2}-\mathbf{{G}}_{i,j-1/2}}{\Delta y}+\bar{\mathbf{S}}_{ij}\right)\,,$ (73)
@@
The non-conservative products are only non-zero for the first equation and are computed using an upwind formulation based on a high-order (7-th order) WENO reconstruction [56, 45]. The diffusive fluxes are only non-zero for the first equation, that is
@@
where $\widetilde{k}=0.5(k_{i+1}+k_{i})$, so it is the mean of the non-linear diffusion in Eq. (28).